40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 18 18:33:57 2014, ericq, Update, Green Locking, Yarm Green PDH yLoop.pdfcrap.pdf
    Reply  Tue Aug 19 01:01:36 2014, Jenne, Rana, Update, Green Locking, Yarm Green PDH YPDH_OLG.pdfYPDH_noise.pdf
       Reply  Tue Aug 19 18:32:40 2014, ericq, Update, Green Locking, Yarm Green PDH 
       Reply  Tue Aug 19 23:11:15 2014, Jenne, Update, Green Locking, Yarm Green PDH ErrAndCtrlSpectra_VoltsPerRtHz.png
       Reply  Wed Aug 20 02:38:41 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement NoiseConsideration.pdf
          Reply  Wed Aug 20 18:05:18 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement YPDH_noise.pdf
       Reply  Wed Aug 20 04:09:21 2014, ericq, Update, Green Locking, Xarm Green PDH Xbodes.pdfXspectra.pdfscopeSweep.jpg
          Reply  Wed Aug 20 16:10:43 2014, ericq, Update, Green Locking, Xarm Green PDH XspectraCombined.pdf
             Reply  Wed Aug 20 21:09:16 2014, ericq, Update, Green Locking, Xarm Green PDH scopeSweep.jpgXspectra.pdfYspectra.pdfXYcomp.pdf
Message ID: 10417     Entry time: Wed Aug 20 21:09:16 2014     In reply to: 10415
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: Xarm Green PDH 

I remeasured all of the noise spectra again today, making sure the input attenuation was as low as it could safely be. I also got a snap of the y green PDH signal; it's fairly larger than I saw the other day, which is good. I used this to calibrate the error signal voltage spectra. 

scopeSweep.jpg

Here are the noise traces for each arm. During these measurements GTRX was about .6, GTRY about 1.0 The Yarm noise doesn't look so good: the error signal is just barely above the mixer+lowpass output noise, and the RMS is plauged by 60Hz lines. (Is this related to what we see in IR TRY sometimes?)

Xspectra.pdfYspectra.pdf

Here are the arms error signals compared directly:

XYcomp.pdf

ELOG V3.1.3-